What is the momentum of a 145 g baseball moving at a speed of 26.5 m / s?


1
Expert's answer
2021-01-07T10:55:07-0500

It's known that the momentum "p" is "mv," so

"p=0.145\cdot 26.5=3.84 ~\frac{kg\cdot m}{s}."